# Homological algebra of pro-Lie Polish abelian groups

### v1

@@ 2.4. Derived functors
Suppose that $\mathcal{A}$ and $\mathcal{R}$ are exact categories, and $F:\mathcal{A}\rightarrow\mathcal{R}$ is a functor. Then $F$ induces a triangulated functor $\mathrm{K}^{+}\left(\mathcal{A}\right)\rightarrow\mathrm{K}^{+}\left(\mathcal{R}\right)$ [31, Definition 2.1.1], which we still denote by $F$. A total right derived functor for $F:\mathrm{K}^{+}\left(\mathcal{A}\right)\rightarrow\mathrm{K}^{+}\left(\mathcal{R}\right)$ [7, Section 10.6]—see also [34, Definition 1.3.1]—is a triangulated functor $\mathrm{R}F:\mathrm{D}^{+}\left(\mathcal{A}\right)\rightarrow\mathrm{D}^{+}\left(\mathcal{R}\right)$ together with a morphism $\mu:Q_{\mathcal{R}}\circ F\Rightarrow\mathrm{R}F\circ Q_{\mathcal{A}}$ of triangulated functors that satisfies the following universal property: for every triangulated functor $G:\mathrm{D}^{+}\left(\mathcal{A}\right)\rightarrow\mathrm{D}^{+}\left(\mathcal{R}\right)$ and morphism $\nu:Q_{\mathcal{R}}\circ F\rightarrow G\circ Q_{\mathcal{A}}$ of triangulated functors, there exists a unique morphism of triangulated functors $\sigma:\mathrm{R}F\Rightarrow G$ such that $\nu=\left(\sigma Q_{\mathcal{A}}\right)\circ\mu$.
If $\mathcal{R}$ is quasi-abelian, then $\mathrm{H}^{0}\circ F$ is a cohomological functor [36, Section 1.1, Definition 3.1]; see also [31, Definition 1.1.7]. A right cohomological derived functor for $\mathrm{H}^{0}\circ F$ is defined in a similar fashion, via a suitable universal property [36, Section 2.2, Definition 1.4].
@@
    generating [24, Definition 8.3.21(v)] if for each object $A$ of $\mathcal{A}$ there exists an admissible epimorphism $C\rightarrow A$ with $C$ in $\mathcal{C}$;
@@
    $0\rightarrow C^{\prime}\rightarrow C\rightarrow A^{\prime\prime}\rightarrow 0$
    in $\mathcal{A}$ with $C$ and $C^{\prime}$ in $\mathcal{C}$, $A^{\prime\prime}$ is isomorphic to an object of $\mathcal{C}$;
@@
We have the following result [7, Theorem 10.22 and Remark 10.23]; see also [24, Proposition 13.2.2], [34, Lemma 1.3.3 and Lemma 1.3.4], and [21, Corollary 3.10].
@@
As a consequence of Lemma 2.6 we have the following; see [7, Section 10.6], [24, Definition 10.3.2, Proposition 10.3.3, Corollary 13.3.8], [34, Proposition 1.3.5], and [21, Proposition 4.10].
@@
This yields a triangulated functor $\mathrm{R}F:\mathrm{D}^{b}\left(\mathcal{A}\right)\rightarrow\mathrm{D}^{b}\left(\mathcal{R}\right)$. Defining $\mu_{A}:=F\left(\eta_{A}\right):F\left(A\right)\rightarrow\left(\mathrm{R}F\right)\left(A\right)$ for each bounded complex $A$ over $\mathcal{A}$ yields an isomorphism $\mu:Q_{\mathcal{R}}F\Rightarrow\left(\mathrm{R}F\right)Q_{\mathcal{A}}$ of triangulated functors. We have that $\left(\mathrm{R}F,\mu\right)$ is the total right derived functor of $F$. Furthermore, if $\mathcal{R}$ is a quasi-abelian category, then $\mathrm{H}^{0}\circ\mathrm{R}F$ is a cohomological right derived functor of $\mathrm{H}^{0}\circ F$.
Suppose that $\mathcal{A}$, $\mathcal{B}$, $\mathcal{R}$ are exact categories, and $F:\mathcal{A}^{\mathrm{op}}\times\mathcal{B}\rightarrow\mathcal{R}$ is a functor. Given bounded complexes $A$ over $\mathcal{A}$ and $B$ over $\mathcal{B}$, we have a corresponding double complex $F\left(A,B\right)$ over $\mathcal{R}$. We let $F^{\bullet}\left(A,B\right)$ be the total complex associated with $F\left(A,B\right)$, which is well-defined and bounded since $A$ and $B$ are bounded and hence $F\left(A,B\right)$ has only finitely many nonzero entries in each diagonal. This defines a triangulated functor $F^{\bullet}:\mathrm{K}^{b}\left(\mathcal{A}\right)^{\mathrm{op}}\times\mathrm{K}^{b}\left(\mathcal{B}\right)\rightarrow\mathrm{K}^{b}\left(\mathcal{R}\right)$; see [24, Proposition 11.6.4]. The same proof as Proposition 2.7 gives the following.
@@
Suppose that $F:\mathcal{A}^{\mathrm{op}}\times\mathcal{B}\rightarrow\mathcal{R}$ is a functor between exact categories, $\mathcal{C}$ is an generating full subcategory of $\mathcal{A}$ closed under subobjects, and $\mathcal{D}$ is a cogenerating full subcategory of $\mathcal{B}$ closed under quotients. Assume that for every object $C$ of $\mathcal{C}$ and $D$ of $\mathcal{D}$, $F\left(C,-\right)$ is exact on $\mathcal{D}$ and $F\left(-,D\right)$ is exact on $\mathcal{C}$. For bounded complexes $A$ and $B$ over $\mathcal{A}$ and $\mathcal{B}$, respectively, pick bounded complexes $C_{A}$ and $D_{B}$ over $\mathcal{C}$ and $\mathcal{D}$, respectively, together with quasi-isomorphism $\eta_{A}:C_{A}\rightarrow A$ and $\eta_{B}:B\rightarrow D_{B}$, and define
@@
This yields a triangulated functor $\mathrm{R}F^{\bullet}:\mathrm{D}^{b}\left(\mathcal{A}\right)^{\mathrm{op}}\times\mathrm{D}^{b}\left(\mathcal{B}\right)\rightarrow\mathrm{D}^{b}\left(\mathcal{R}\right)$. Defining $\mu_{A}:=F\left(\eta_{A},\eta_{B}\right):F^{\bullet}\left(A,B\right)\rightarrow\left(\mathrm{R}F\right)\left(A\right)$ for each bounded complex $A$ over $\mathcal{A}$ yields an isomorphism $\mu:Q_{\mathcal{R}}F^{\bullet}\Rightarrow\left(\mathrm{R}F\right)Q_{\mathcal{A}}$ of triangulated functors. We have that $\left(\mathrm{R}F,\mu\right)$ is the total right derived functor of $F^{\bullet}$. Furthermore, if $\mathcal{R}$ is a quasi-abelian category, then $\mathrm{H}^{0}\circ\mathrm{R}F^{\bullet}$ is a cohomological right derived functor of $\mathrm{H}^{0}\circ F^{\bullet}$.
@@ 2.5. Injective and projective objects
For objects $A$ and $B$ of $\mathcal{A}$ and $n\in\mathbb{Z}$, one defines $\mathrm{Ext}^{n}\left(A,B\right):=\mathrm{Hom}_{\mathrm{D}\left(\mathcal{A}\right)}\left(A,B[n]\right)$. One says that $\mathcal{A}$ has finite homological dimension if there exists $d\in\mathbb{Z}$ such that $\mathrm{Ext}^{n}\left(A,B\right)=0$ for all $n>d$ and objects $A$ and $B$ of $\mathcal{A}$. In this case, the least such $d$ is called the homological dimension $\mathrm{hd}\left(\mathcal{A}\right)$ of $\mathcal{A}$ [24, Exercise 13.8]. If $\mathcal{A}$ has enough injectives and $d\geq 0$, then $\mathrm{hd}\left(\mathcal{A}\right)\leq d$ if and only if for every exact sequence $X^{0}\rightarrow X^{1}\rightarrow\cdots\rightarrow X^{d}\rightarrow 0$ in $\mathcal{A}$ with $X^{0},X^{1},\ldots,X^{d-1}$ injective, we have that $X^{d}$ is also injective [24, Exercise 13.8]. In this case, every object of $\mathcal{A}$ has an injective resolution $I_{A}$ with $I_{A}^{k}=0$ for $k\in\mathbb{Z}\setminus\left\{0,1,\ldots,d\right\}$.
If $\mathcal{A}$ has enough injectives and finite homological dimension, then for every (bounded) complex $A$ over $\mathcal{A}$ there exists a quasi-isomorphism $\mu_{A}:A\rightarrow I_{A}$ where $I_{A}$ is a (bounded) complex over $\mathcal{I}$ [22, Corollary I.7.7]. Furthermore, the inclusion $\mathcal{I}\rightarrow\mathcal{A}$ induces an equivalence of categories $\mathrm{K}\left(\mathcal{I}\right)\rightarrow\mathrm{D}\left(\mathcal{A}\right)$, which restricts to an equivalence of categories $\mathrm{K}^{b}\left(\mathcal{I}\right)\rightarrow\mathrm{D}^{b}\left(\mathcal{A}\right)$ [22, Proposition IX.2.12]; see also [24, Proposition 13.2.2].
@@
Let $\mathcal{A}$ and $\mathcal{R}$ be exact categories. Suppose that $\mathcal{A}$ has enough injectives or enough projectives, and homological dimension at most $1$. If $F:\mathcal{A}^{\mathrm{op}}\times\mathcal{A}\rightarrow\mathcal{R}$ is a functor, then $F^{\bullet}:\mathrm{K}^{b}\left(\mathcal{A}\right)^{\mathrm{op}}\times\mathrm{K}^{b}\left(\mathcal{A}\right)\rightarrow\mathrm{K}^{b}\left(\mathcal{R}\right)$ has a total right derived functor $\mathrm{R}F:\mathrm{D}^{b}\left(\mathcal{A}\right)^{\mathrm{op}}\times\mathrm{D}^{b}\left(\mathcal{A}\right)\rightarrow\mathrm{D}^{b}\left(\mathcal{R}\right)$. If $\mathcal{R}$ is quasi-abelian, then $\mathrm{H}^{0}\circ\mathrm{R}F$ is a cohomological derived functor of $\mathrm{H}^{0}\circ F^{\bullet}$.
For the purposes of the following corollary, we say that a triangulated category $\mathcal{T}$ is enriched over a quasi-abelian category $\mathcal{M}$ if for objects $A$ and $B$ of $\mathcal{T}$, $\mathrm{Hom}\left(A,-\right)$ and $\mathrm{Hom}\left(-,B\right)$ are functors to $\mathcal{M}$ such that $\mathrm{Hom}\left(A,-\right)$ is a homological functor in the sense of [31, Definition 1.1.7] when regarded as a functor $\mathcal{T}\rightarrow\mathrm{LH}\left(\mathcal{M}\right)$, and $\mathrm{Hom}\left(-,B\right)$ is a cohomological functor when regarded as a functor $\mathcal{T}^{\mathrm{op}}\rightarrow\mathrm{LH}\left(\mathcal{M}\right)$. As a particular instance of Proposition 2.8, one has the following.
@@
Suppose that $\mathcal{M}$ and $\mathcal{A}$ are quasi-abelian categories, where $\mathcal{A}$ is $\mathcal{M}$-enriched. Assume that $\mathcal{A}$ has enough injectives or enough projectives, and homological dimension at most $1$. Then $\mathrm{Hom}^{\bullet}:\mathrm{K}^{b}\left(\mathcal{A}\right)^{\mathrm{op}}\times\mathrm{K}^{b}\left(\mathcal{A}\right)\rightarrow\mathrm{K}^{b}\left(\mathcal{M}\right)$ has a total right derived functor $\mathrm{RHom}:\mathrm{D}^{b}\left(\mathcal{A}\right)^{\mathrm{op}}\times\mathrm{D}^{b}\left(\mathcal{A}\right)\rightarrow\mathrm{D}^{b}\left(\mathcal{M}\right)$, and $\mathrm{Ext}^{0}:=\mathrm{H}^{0}\circ\mathrm{RHom}$ is a cohomological derived functor of $\mathrm{Hom}_{\mathrm{K}\left(\mathcal{A}\right)}:\mathrm{K}^{b}\left(\mathcal{A}\right)^{\mathrm{op}}\times\mathrm{K}^{b}\left(\mathcal{A}\right)\rightarrow\mathrm{LH}(\mathcal{M)}$. Furthermore, for bounded complexes $A$ and $B$ over $\mathcal{A}$, $\mathrm{Ext}^{0}\left(A,B\right)$ is naturally isomorphic as an abelian group to $\mathrm{Hom}_{\mathrm{D}\left(\mathcal{A}\right)}\left(A,B\right)$. This isomorphism turns $\mathrm{D}^{b}\left(\mathcal{A}\right)$ into a triangulated category enriched over $\mathrm{LH}\left(\mathcal{M}\right)$.
@@
For the second assertion, observe that by [24, Proposition 11.7.3], we have that $\mathrm{Hom}_{\mathrm{K}\left(\mathcal{A}\right)}$ is naturally isomorphic to $\mathrm{H}^{0}\circ\mathrm{Hom}^{\bullet}$. Therefore, by the first assertion we have that $\mathrm{RHom}$ is a total right derived functor of $\mathrm{Hom}^{\bullet}$, and $\mathrm{Ext}^{0}:=\mathrm{H}^{0}\circ\mathrm{RHom}$ is a cohomological right derived functor of $\mathrm{Hom}_{\mathrm{K}\left(\mathcal{A}\right)}\cong\mathrm{H}^{0}\circ\mathrm{Hom}^{\bullet}$. Furthermore, we have that $\mathrm{Ext}^{0}$ is naturally isomorphic to $\mathrm{Hom}_{\mathrm{D}^{b}\left(\mathcal{A}\right)}$ by [24, Proposition 11.7.3]. This allows one to regard $\mathrm{Hom}_{\mathrm{D}^{b}\left(\mathcal{A}\right)}$ as a functor to $\mathrm{D}^{b}\left(\mathcal{A}\right)^{\mathrm{op}}\times\mathrm{D}^{b}\left(\mathcal{A}\right)\rightarrow\mathrm{LH}\left(\mathcal{M}\right)$, which turns $\mathrm{D}^{b}\left(\mathcal{A}\right)$ into a triangulated category enriched over $\mathrm{LH}\left(\mathcal{M}\right)$. ∎

### v2

@@ 2.4. Derived functors
Suppose that $\mathcal{A}$ and $\mathcal{R}$ are exact categories, and $F:\mathcal{A}\rightarrow\mathcal{R}$ is a functor. Then $F$ induces a triangulated functor $\mathrm{K}^{+}\left(\mathcal{A}\right)\rightarrow\mathrm{K}^{+}\left(\mathcal{R}\right)$ [32, Definition 2.1.1], which we still denote by $F$. A total right derived functor for $F:\mathrm{K}^{+}\left(\mathcal{A}\right)\rightarrow\mathrm{K}^{+}\left(\mathcal{R}\right)$ [7, Section 10.6]—see also [35, Definition 1.3.1]—is a triangulated functor $\mathrm{R}F:\mathrm{D}^{+}\left(\mathcal{A}\right)\rightarrow\mathrm{D}^{+}\left(\mathcal{R}\right)$ together with a morphism $\mu:Q_{\mathcal{R}}\circ F\Rightarrow\mathrm{R}F\circ Q_{\mathcal{A}}$ of triangulated functors that satisfies the following universal property: for every triangulated functor $G:\mathrm{D}^{+}\left(\mathcal{A}\right)\rightarrow\mathrm{D}^{+}\left(\mathcal{R}\right)$ and morphism $\nu:Q_{\mathcal{R}}\circ F\rightarrow G\circ Q_{\mathcal{A}}$ of triangulated functors, there exists a unique morphism of triangulated functors $\sigma:\mathrm{R}F\Rightarrow G$ such that $\nu=\left(\sigma Q_{\mathcal{A}}\right)\circ\mu$.
If $\mathcal{R}$ is quasi-abelian, then $\mathrm{H}^{0}\circ F$ is a cohomological functor [37, Section 1.1, Definition 3.1]; see also [32, Definition 1.1.7]. A right cohomological derived functor for $\mathrm{H}^{0}\circ F$ is defined in a similar fashion, via a suitable universal property [37, Section 2.2, Definition 1.4].
@@
    generating [25, Definition 8.3.21(v)] if for each object $A$ of $\mathcal{A}$ there exists an admissible epimorphism $C\rightarrow A$ with $C$ in $\mathcal{C}$;
@@
    $0\rightarrow A^{\prime}\rightarrow C\rightarrow A^{\prime\prime}\rightarrow 0$
    in $\mathcal{A}$ with $C$ in $\mathcal{C}$, $A^{\prime\prime}$ is isomorphic to an object of $\mathcal{C}$;
@@
We have the following result [25, Proposition 13.2.2(b)]; see also [7, Theorem 10.22 and Remark 10.23], [25, Proposition 13.2.2], [35, Lemma 1.3.3 and Lemma 1.3.4], and [22, Corollary 3.10].
@@
As a consequence of Lemma 2.10 we have the following; see [7, Section 10.6], [25, Definition 10.3.2, Proposition 10.3.3, Corollary 13.3.8], [35, Proposition 1.3.5], and [22, Remark 4.10].
@@
This yields a triangulated functor $\mathrm{R}F:\mathrm{D}^{b}\left(\mathcal{A}\right)\rightarrow\mathrm{D}^{b}\left(\mathcal{R}\right)$. Defining $\mu_{A}:=F\left(\eta_{A}\right):F\left(A\right)\rightarrow\left(\mathrm{R}F\right)\left(A\right)$ for each bounded complex $A$ over $\mathcal{A}$ yields a morphism $\mu:Q_{\mathcal{R}}F\Rightarrow\left(\mathrm{R}F\right)Q_{\mathcal{A}}$ of triangulated functors. We have that $\left(\mathrm{R}F,\mu\right)$ is the total right derived functor of $F$. Furthermore, if $\mathcal{R}$ is a quasi-abelian category, then $\mathrm{H}^{0}\circ\mathrm{R}F$ is a cohomological right derived functor of $\mathrm{H}^{0}\circ F$.
Suppose that $\mathcal{A}$, $\mathcal{B}$, $\mathcal{R}$ are exact categories, and $F:\mathcal{A}^{\mathrm{op}}\times\mathcal{B}\rightarrow\mathcal{R}$ is a functor. Given bounded complexes $A$ over $\mathcal{A}$ and $B$ over $\mathcal{B}$, we have a corresponding double complex $F\left(A,B\right)$ over $\mathcal{R}$. We let $F^{\bullet}\left(A,B\right)$ be the total complex associated with $F\left(A,B\right)$, which is well-defined and bounded since $A$ and $B$ are bounded and hence $F\left(A,B\right)$ has only finitely many nonzero entries in each diagonal. This defines a triangulated functor $F^{\bullet}:\mathrm{K}^{b}\left(\mathcal{A}\right)^{\mathrm{op}}\times\mathrm{K}^{b}\left(\mathcal{B}\right)\rightarrow\mathrm{K}^{b}\left(\mathcal{R}\right)$; see [25, Proposition 11.6.4]. The same proof as Proposition 2.11 gives the following.
@@
Suppose that $F:\mathcal{A}^{\mathrm{op}}\times\mathcal{B}\rightarrow\mathcal{R}$ is a functor between exact categories, $\mathcal{C}$ is an generating full subcategory of $\mathcal{A}$ closed under subobjects, and $\mathcal{D}$ is a cogenerating full subcategory of $\mathcal{B}$ closed under quotients. Assume that for every object $C$ of $\mathcal{C}$ and $D$ of $\mathcal{D}$, $F\left(C,-\right)$ is exact on $\mathcal{D}$ and $F\left(-,D\right)$ is exact on $\mathcal{C}$. For bounded complexes $A$ and $B$ over $\mathcal{A}$ and $\mathcal{B}$, respectively, pick bounded complexes $C_{A}$ and $D_{B}$ over $\mathcal{C}$ and $\mathcal{D}$, respectively, together with quasi-isomorphisms $\eta_{A}:C_{A}\rightarrow A$ and $\eta_{B}:B\rightarrow D_{B}$, and define
@@
This yields a triangulated functor $\mathrm{R}F^{\bullet}:\mathrm{D}^{b}\left(\mathcal{A}\right)^{\mathrm{op}}\times\mathrm{D}^{b}\left(\mathcal{B}\right)\rightarrow\mathrm{D}^{b}\left(\mathcal{R}\right)$. Defining $\mu_{A}:=F\left(\eta_{A},\eta_{B}\right):F^{\bullet}\left(A,B\right)\rightarrow\left(\mathrm{R}F\right)\left(A\right)$ for each bounded complex $A$ over $\mathcal{A}$ yields a morphism $\mu:Q_{\mathcal{R}}F^{\bullet}\Rightarrow\left(\mathrm{R}F\right)\left(Q_{\mathcal{A}}\times Q_{\mathcal{B}}\right)$ of triangulated functors. We have that $\left(\mathrm{R}F,\mu\right)$ is the total right derived functor of $F^{\bullet}$. Furthermore, if $\mathcal{R}$ is a quasi-abelian category, then $\mathrm{H}^{0}\circ\mathrm{R}F^{\bullet}$ is a cohomological right derived functor of $\mathrm{H}^{0}\circ F^{\bullet}$.
@@ 2.5. Injective and projective objects
For objects $A$ and $B$ of $\mathcal{A}$ and $n\in\mathbb{Z}$, one defines $\mathrm{Ext}^{n}\left(A,B\right):=\mathrm{Hom}_{\mathrm{D}\left(\mathcal{A}\right)}\left(A,B[n]\right)$. One says that $\mathcal{A}$ has finite homological dimension if there exists $d\in\mathbb{Z}$ such that $\mathrm{Ext}^{n}\left(A,B\right)=0$ for all $n>d$ and objects $A$ and $B$ of $\mathcal{A}$. In this case, the least such $d$ is called the homological dimension $\mathrm{hd}\left(\mathcal{A}\right)$ of $\mathcal{A}$ [25, Exercise 13.8]. If $\mathcal{A}$ has enough injectives and $d\geq 0$, then $\mathrm{hd}\left(\mathcal{A}\right)\leq d$ if and only if for every exact sequence $X^{0}\rightarrow X^{1}\rightarrow\cdots\rightarrow X^{d}\rightarrow 0$ in $\mathcal{A}$ with $X^{0},X^{1},\ldots,X^{d-1}$ injective, we have that $X^{d}$ is also injective [25, Exercise 13.8]. In this case, every object of $\mathcal{A}$ has an injective resolution $I_{A}$ with $I_{A}^{k}=0$ for $k\in\mathbb{Z}\setminus\left\{0,1,\ldots,d\right\}$.
If $\mathcal{A}$ has enough injectives and finite homological dimension, then for every (bounded) complex $A$ over $\mathcal{A}$ there exists a quasi-isomorphism $\mu_{A}:A\rightarrow I_{A}$ where $I_{A}$ is a (bounded) complex over $\mathcal{I}$ [23, Corollary I.7.7]. Furthermore, the inclusion $\mathcal{I}\rightarrow\mathcal{A}$ induces an equivalence of categories $\mathrm{K}\left(\mathcal{I}\right)\rightarrow\mathrm{D}\left(\mathcal{A}\right)$, which restricts to an equivalence of categories $\mathrm{K}^{b}\left(\mathcal{I}\right)\rightarrow\mathrm{D}^{b}\left(\mathcal{A}\right)$ [23, Proposition IX.2.12]; see also [25, Proposition 13.2.2]. As a consequence of Proposition 2.12 one obtains the following:
@@
Let $\mathcal{A}$ and $\mathcal{R}$ be exact categories. Let $F:\mathcal{A}^{\mathrm{op}}\times\mathcal{A}\rightarrow\mathcal{R}$ be a functor. Suppose that either:
  • $\mathcal{A}$ has enough injectives, and $F\left(-,I\right):\mathcal{A}^{\mathrm{op}}\rightarrow\mathcal{R}$ is exact for every injective object $I$ of $\mathcal{A}$, or
  • $\mathcal{A}$ has enough projectives, and $F\left(P,-\right):\mathcal{A}^{\mathrm{op}}\rightarrow\mathcal{R}$ is exact for every projective object $P$ of $\mathcal{A}$.
Suppose furthermore that $\mathcal{A}$ has homological dimension at most $1$. Then $F^{\bullet}:\mathrm{K}^{b}\left(\mathcal{A}\right)^{\mathrm{op}}\times\mathrm{K}^{b}\left(\mathcal{A}\right)\rightarrow\mathrm{K}^{b}\left(\mathcal{R}\right)$ has a total right derived functor $\mathrm{R}F:\mathrm{D}^{b}\left(\mathcal{A}\right)^{\mathrm{op}}\times\mathrm{D}^{b}\left(\mathcal{A}\right)\rightarrow\mathrm{D}^{b}\left(\mathcal{R}\right)$. If $\mathcal{R}$ is quasi-abelian, then $\mathrm{H}^{0}\circ\mathrm{R}F$ is a cohomological derived functor of $\mathrm{H}^{0}\circ F^{\bullet}$.
As a particular instance of Proposition 2.12, one has the following.
@@
Suppose that $\mathcal{M}$ is a quasi-abelian category of abelian groups, and $\mathcal{A}$ is a quasi-abelian $\mathcal{M}$-category. Assume that $\mathcal{A}$ has enough injectives or enough projectives, and homological dimension at most $1$. Then $\mathrm{Hom}^{\bullet}:\mathrm{K}^{b}\left(\mathcal{A}\right)^{\mathrm{op}}\times\mathrm{K}^{b}\left(\mathcal{A}\right)\rightarrow\mathrm{K}^{b}\left(\mathcal{M}\right)$ has a total right derived functor $\mathrm{RHom}:\mathrm{D}^{b}\left(\mathcal{A}\right)^{\mathrm{op}}\times\mathrm{D}^{b}\left(\mathcal{A}\right)\rightarrow\mathrm{D}^{b}\left(\mathcal{M}\right)$, and $\mathrm{Ext}^{0}:=\mathrm{H}^{0}\circ\mathrm{RHom}$ is a cohomological derived functor of $\mathrm{Hom}_{\mathrm{K}\left(\mathcal{A}\right)}:\mathrm{K}^{b}\left(\mathcal{A}\right)^{\mathrm{op}}\times\mathrm{K}^{b}\left(\mathcal{A}\right)\rightarrow\mathrm{LH}(\mathcal{M)}$. Furthermore, for bounded complexes $A$ and $B$ over $\mathcal{A}$, $\mathrm{Ext}^{0}\left(A,B\right)$ is naturally isomorphic as an abelian group to $\mathrm{Hom}_{\mathrm{D}\left(\mathcal{A}\right)}\left(A,B\right)$. This isomorphism turns $\mathrm{D}^{b}\left(\mathcal{A}\right)$ into a triangulated category enriched over $\mathrm{LH}\left(\mathcal{M}\right)$.
@@
For the second assertion, observe that by [25, Proposition 11.7.3], we have that $\mathrm{Hom}_{\mathrm{K}\left(\mathcal{A}\right)}$ is naturally isomorphic to $\mathrm{H}^{0}\circ\mathrm{Hom}^{\bullet}$. Therefore, by the first assertion we have that $\mathrm{RHom}$ is a total right derived functor of $\mathrm{Hom}^{\bullet}$, and $\mathrm{Ext}^{0}:=\mathrm{H}^{0}\circ\mathrm{RHom}$ is a cohomological right derived functor of $\mathrm{Hom}_{\mathrm{K}\left(\mathcal{A}\right)}\cong\mathrm{H}^{0}\circ\mathrm{Hom}^{\bullet}$. Furthermore, we have that $\mathrm{Ext}^{0}$ is naturally isomorphic to $\mathrm{Hom}_{\mathrm{D}^{b}\left(\mathcal{A}\right)}$ by [25, Proposition 11.7.3]. This allows one to regard $\mathrm{Hom}_{\mathrm{D}^{b}\left(\mathcal{A}\right)}$ as a functor to $\mathrm{D}^{b}\left(\mathcal{A}\right)^{\mathrm{op}}\times\mathrm{D}^{b}\left(\mathcal{A}\right)\rightarrow\mathrm{LH}\left(\mathcal{M}\right)$, which turns $\mathrm{D}^{b}\left(\mathcal{A}\right)$ into a triangulated category enriched over $\mathrm{LH}\left(\mathcal{M}\right)$. ∎